I need help with this question?
Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 1/2 hours. How quickly can all three fill the pool?

Answer by stanbon(75887)

Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 1/2 hours. How quickly can all three fill the pool?
------
Jim rate: 1/(1/2) = 2 jobs/hr
---
Sue rate: (1/(3/4)) = (4/3) jobs/hr
---
Tony rate = 1/(3/2) = 2/3 job/hr
-------------------------------------
Together rate: 1/x jobs/hr
----------------------------------
Equation:
rate + rate + rate = together rate
2 + 4/3 + 2/3 = 1/x
----
Multiply thru by 3x to get:
6x + 4x + 2x = 3
12x = 3
x = 1/4 hr.
==================
Cheers,
Stan H.
=========